A plate clutch has two disA plate clutch has two discs on driven shaft, each effective on both sides, and of outer and inner radii 280 mm and 140 mm, respectively. It transmits 45 kW at 2400 rev/min and μ = 0.32 for the friction material. Calculate: 4.1 The axial force needed assuming uniform pressure. (7) 4.2 The force that would be needed assuming uniform wear. (5)cs on driven shaft, each effective on both sides, and of outer and inner radii 280 mm and 140 mm, respectively. It transmits 45 kW at 2400 rev/min and μ = 0.32 for the friction material. Calculate: 4.1 The axial force needed assuming uniform pressure. (7) 4.2 The force that would be needed assuming uniform wear. (5)


1
Expert's answer
2021-08-19T14:18:34-0400

1)

"P=\frac At=\frac{\mu F_1 C}t=\mu F_12\pi(R-r)2\pi \nu=4\pi^2\mu F_1\nu(R-r),\\
F_1=\frac P{4\pi^2\mu \nu(R-r)}=636~N,"

2)

"F_2=F_1\frac Rr=1272~N."